An electronic technician wishes to construct a parallel-plate capacitor using rutile (k = 100) as the dielectric. The area of the plates is 1.00cm^(2) . What is the capacitance if the rutile thickness is 1.00mm?
1
Expert's answer
2021-02-17T15:23:04-0500

The capacitance of the capacitor filled with dielectric between its plates can be calculated from the formula:


"C=k\epsilon_0\dfrac{A}{d},""C=100\cdot8.85\cdot10^{-12}\ \dfrac{F}{m}\cdot\dfrac{1.0\ cm^2\cdot(\dfrac{1\ m}{100\ cm})^2}{1.0\cdot10^{-3}\ m}=88.5\ pF."